# Read Between the Hyperplanes: On Spectral Projection and Sampling

### v2

@@ 2 Pairwise Comparisons for Binary Classification
We consider a consistent system of inequalities $Ax^{*}\leq b$ for some matrix $A$ and an unknown intended solution vector $x^{*}$ such that
@@ 3 Clustering
Another central question we investigated concerns the extraction of a compact coreset from an over-determined linear system. We conjecture that any system of size $m\texttimes n$ admits a coreset of cardinality O(cn)O(c*n), where $c$ is a small constant independent of $m$. Identifying such a subset would enable existing algorithms to operate on a substantially reduced system while preserving the original problem’s solution quality. Our approach is to partition the matrix $A$ into clusters of similar rows, treating each cluster as an approximately redundant group whose representative captures the essential structure of that subset. This clustering strategy allows for parallel processing and lowers the overall computational cost without significant loss of accuracy. Throughout our experiments, we use the Kaczmarz algorithm as the baseline solver for evaluating convergence speed and approximation fidelity. The following section elaborates on the theoretical intuition and presents the algorithms that implement this idea.
@@ 3.1 Existence of a “Good” Cluster
The results in Figure 2 confirm that compact coresets can retain the fundamental subspace geometry of the original system even under extreme ill-conditioning. As the coreset factor $c$ grows, the relative error between $x_{B}$ and $x_{A}$ decreases approximately monotonically, suggesting that the selected rows form an increasingly accurate projection of the column space of $A$. The near-constant residual magnitudes across condition numbers indicate that the coreset preserves the dominant left-singular subspace, implying that $B^{\top}B\approx A^{\top}A$ in spectral norm up to a small perturbation. These empirical results align with the intuition that low-rank or leverage-based clustering implicitly enforces subspace embedding guarantees, thereby maintaining both conditioning and numerical stability while drastically reducing the system size.
@@ 3.2.2 Online cluster updating
Another idea is essentially taken from a well-known K-means algorithm. Since we know that the Karzmarz algorithm will eventually converge to a solution, we were trying to find an optimization that will be updated during the iterations. Each fixed number of iterations we were trying to reduce the matrix $A$ that we were working with. Particularly, for an $m\times n$ matrix:
@@
To obtain a more discriminative setting, we therefore evaluated all methods on a synthetically generated ill-conditioned matrix with fixed condition number on the order of $10^{7}$. In Figure 3 we report the convergence of four stochastic Kaczmarz Motzkin (SKM) variants over $2000$ iterations on a system of size $(m,n)=(2000,20)$. The “reduced matrix” (proof-of-concept) variant uses a constant $c=5$, corresponding to a coreset of about $5\%$ of the rows. In this experiment, the reduced matrix method achieves accuracy and approximation error comparable to the full-matrix Hadamard SKM baseline, demonstrating that a suitably selected subset of rows can solve the linear system nearly as well as the original system. However, none of the modified algorithms matches the performance of the original Hadamard SKM, indicating a nontrivial trade-off between reduction and convergence speed in the ill-conditioned regime.
@@ 4.2 Approach
where each $\vec{v}_{j}$ is the $j^{th}$ singular vector of $A$. In the context of our problem, it should not be assumed that knowledge of the singular vectors be known; otherwise, the solution could be reconstructed and the problem would be trivial. However, we analyze the convergence using this knowledge to inspire a new algorithm. Considering that $A=U\Sigma V^{\mathsf{T}}$, we obtain each $c^{(i)}_{n}$ by:
@@ 4.3 Numerical Experimentation
Although this method requires knowledge of the singular value decomposition of $A$, the trend it unveils should be used to inspire methods that improve existing RK algorithms for ill-conditioned systems. We propose some solutions that leverage these properties in Section 5.
@@ 5 Future Work
We remind the reader that the weighted sampling method proposed in Section 4 requires the knowledge of the SVD of the matrix beforehand. Since the process of obtaining the SVD is, itself, expensive and would render the problem trivial, we suggest approximating the singular vectors $v_{j}$ of $A$ iteratively. We suggest two possible approaches: (i) iteratively adapt sampling weights as the singular vectors are approximated, or (ii) perform Average Block Kaczmarz (ABK) [1] with $\beta$ many rows and adjust the projection weight of each row proportionally to the magnitude of its component $c^{(i)}_{n}$ relative to the other rows’.

### v3

# arXiv:2511.03055v3[math.NA] 28 Feb 2026

@@ 2 Pairwise Comparisons for Binary Classification
We consider a consistent system of inequalities $Ax^{*}\leq b$ for some matrix AA and an unknown intended solution vector $x^{*}$ such that
@@ 3 Clustering
Another central question we investigated concerns the extraction of a compact coreset from an over-determined linear system. We conjecture that any system of size $m\texttimes n$ admits a coreset of cardinality O(cn)O(c*n), where $c$ is a small constant independent of $m$. Identifying such a subset would enable existing algorithms to operate on a substantially reduced system while preserving the original problem’s solution quality. Our approach is to partition the matrix AA into clusters of similar rows, treating each cluster as an approximately redundant group whose representative captures the essential structure of that subset. This clustering strategy allows for parallel processing and lowers the overall computational cost without significant loss of accuracy. Throughout our experiments, we use the Kaczmarz algorithm as the baseline solver for evaluating convergence speed and approximation fidelity. The following section elaborates on the theoretical intuition and presents the algorithms that implement this idea.
@@ 3.1 Existence of a “Good” Cluster
The results in Figure 2 confirm that compact coresets can retain the fundamental subspace geometry of the original system even under extreme ill-conditioning. As the coreset factor $c$ grows, the relative error between $x_{B}$ and $x_{A}$ decreases approximately monotonically, suggesting that the selected rows form an increasingly accurate projection of the column space of AA. The near-constant residual magnitudes across condition numbers indicate that the coreset preserves the dominant left-singular subspace, implying that $B^{\top}B\approx A^{\top}A$ in spectral norm up to a small perturbation. These empirical results align with the intuition that low-rank or leverage-based clustering implicitly enforces subspace embedding guarantees, thereby maintaining both conditioning and numerical stability while drastically reducing the system size.
@@ 3.2.2 Online cluster updating
Another idea is essentially taken from a well-known K-means algorithm. Since we know that the Karzmarz algorithm will eventually converge to a solution, we were trying to find an optimization that will be updated during the iterations. Each fixed number of iterations we were trying to reduce the matrix AA that we were working with. Particularly, for an $m\times n$ matrix:
@@
To obtain a more discriminative setting, we therefore evaluated all methods on a synthetically generated ill-conditioned matrix with fixed condition number on the order of $10^{7}$. In Figure 3 we report the convergence of four stochastic Kaczmarz Motzkin (SKM) variants over 20002000 iterations on a system of size $(m,n)=(2000,20)$. The “reduced matrix” (proof-of-concept) variant uses a constant $c=5$, corresponding to a coreset of about $5\%$ of the rows. In this experiment, the reduced matrix method achieves accuracy and approximation error comparable to the full-matrix Hadamard SKM baseline, demonstrating that a suitably selected subset of rows can solve the linear system nearly as well as the original system. However, none of the modified algorithms matches the performance of the original Hadamard SKM, indicating a nontrivial trade-off between reduction and convergence speed in the ill-conditioned regime.
@@ 4.2 Approach
where each $\vec{v}_{j}$ is the $j^{th}$ singular vector of AA. In the context of our problem, it should not be assumed that knowledge of the singular vectors be known; otherwise, the solution could be reconstructed and the problem would be trivial. However, we analyze the convergence using this knowledge to inspire a new algorithm. Considering that $A=U\Sigma V^{\mathsf{T}}$, we obtain each $c^{(i)}_{n}$ by:
@@ 4.3 Numerical Experimentation
Although this method requires knowledge of the singular value decomposition of AA, the trend it unveils should be used to inspire methods that improve existing RK algorithms for ill-conditioned systems. We propose some solutions that leverage these properties in Section 5.
@@ 5 Future Work
We remind the reader that the weighted sampling method proposed in Section 4 requires the knowledge of the SVD of the matrix beforehand. Since the process of obtaining the SVD is, itself, expensive and would render the problem trivial, we suggest approximating the singular vectors $v_{j}$ of AA iteratively. We suggest two possible approaches: (i) iteratively adapt sampling weights as the singular vectors are approximated, or (ii) perform Average Block Kaczmarz (ABK) [1] with $\beta$ many rows and adjust the projection weight of each row proportionally to the magnitude of its component $c^{(i)}_{n}$ relative to the other rows’.